# Coupling without Communication and Drafter-Invariant Speculative Decoding

### v2

@@ Abstract
Suppose Alice has a distribution $\mathcal{P}$ and Bob has a distribution $\mathcal{Q}$. Alice wants to generate a sample $a\sim\mathcal{P}$ and Bob a sample $b\sim\mathcal{Q}$ such that $a=b$ with has as high of probability as possible. It is well-known that, by sampling from an optimal coupling between the distributions, Alice and Bob can achieve $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$, where $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ is the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$.
What if Alice and Bob must solve this same problem without communicating at all? Perhaps surprisingly, with access to public randomness, they can still achieve $\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}\geq 1-2D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. In fact, this bound can be obtained using a simple protocol based on the Weighted MinHash algorithm.
In this work, we explore the “communication-free coupling” problem in greater depth. First, we show that an equally simple protocol based on “Gumbel sampling” matches the worst-case guarantees of the Weighted MinHash approach, but tends to perform better in practice. Conversely, we prove that both approaches are actually sharp: no communication-free protocol can achieve $\Pr[a=b]>\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ in the worst-case. Finally, we prove that, for distributions over $n$ items, there exists a scheme that uses just $O(\log(n/\epsilon))$ bits of communication to achieve $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-\epsilon$, i.e. to essentially match optimal coupling.
Beyond our theoretical results, we demonstrate an application of “communication-free coupling” to speculative decoding, a recent method for accelerating autoregressive large language models [Leviathan, Kalman, Matias, ICML 2023]. We show that communication-free protocols yield a variant of speculative decoding that we call Drafter-Invariant Speculative Decoding, which has the desirable property that the output of the method is fixed given a fixed random seed, regardless of what drafter is used for speculation. Codes are available at https://github.com/majid-daliri/DISD
@@ 1 Introduction
The goal of this paper is to introduce a problem, “communication-free coupling”, that sounds difficult at first glance, but is actually easy to solve in an optimal way. Suppose we have two cooperating parties, Alice and Bob. Alice has a distribution $\mathcal{P}$ and Bob has a distribution $\mathcal{Q}$, which, for simplicity, we assume are discrete distributions over the set $\{1,\ldots,n\}$. The goal is for Alice to sample $a\in\{1,\ldots,n\}$ according to $\mathcal{P}$ and Bob to sample $b\in\{1,\ldots,n\}$ according to $\mathcal{Q}$, in such a way that we maximize the probability that $a=b$. We are interested in how well Alice and Bob can solve this problem without communicating with each other. We do allow them access to a source of shared random numbers.
@@
where $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ is the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$ (Wu, 2020, Section 4.3). To do so, they construct an optimal coupling between the distributions, i.e., a distribution $\mathcal{C}$ over $\{1,\ldots,n\}\times\{1,\ldots,n\}$ with marginals $\mathcal{P}$ and $\mathcal{Q}$ that maximizes $\Pr_{(a,b)\sim\mathcal{C}}[a=b]$. In fact, the total variation distance is defined as $\max_{\text{couplings }\mathcal{C}}\Pr_{(a,b)\sim\mathcal{C}}[a\neq b]$, so the bound in (1) is optimal, with or without communication.
While an optimal coupling between $\mathcal{P}$ and $\mathcal{Q}$ has a simple closed form (see Section 2 for details), sampling from the coupling requires knowledge of both $\mathcal{P}$ and $\mathcal{Q}$. I.e., sampling from an optimal coupling requires communication between Alice and Bob. However, it turns out that it is possible to sample from a near-optimal coupling without any communication at all. In particular, the starting point of our work is a straightforward observation, that is implicit in prior work (Manasse et al., 2010):
@@
There is a communication-free protocol by which Alice samples $a\sim\mathcal{P}$ and Bob $b\sim\mathcal{Q}$ such that:
@@
Since $0\leq D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})\leq 1$, (2) is always larger than $1-2D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. We conclude that it is possible to nearly match the optimum in (1) without communication. 1 follows from an analysis of the popular “Weighted MinHash” method (Manasse et al., 2010; Ioffe, 2010; Shrivastava, 2016; Christiani, 2020). We provide a sharp analysis of the method (with an exact expression for $\Pr[a=b]$) in Section 3, 5.
@@ 1.1 Motivation
The main theoretical question we address in this work is whether the bound of 1 is optimal, or if it possible to get closer to the ideal $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ bound. To cut to the point: no, it is not. 1 is optimal in the worst-case. Before discussing our results, however, we provide some motivation for studying communication-free coupling. It turns out that this problem has a direct application to a technique called speculative decoding that has gained traction for its ability to accelerate inference in autoregressive language models (Leviathan et al., 2023; Chen et al., 2023; Sun et al., 2023), one of the most important algorithmic challenges in modern machine learning (Hoffmann et al., 2022; Zhou et al., 2024).
@@
However, speculate decoding suffers from one major and subtle drawback: if the small “drafter” neural network changes (i.e., because of an update to improve the model), the use of an optimal coupling means that tokens generated by the state-of-the-art neural network will change. This can be potentially problematic in applications of autoregressive language models, where the expected and desired behavior is that the output of a model is fixed given a fixed random seed. It allows for researchers and other users to reliably reproduce results and leads to easier unit testing and debugging.
@@ 1.2 Our Contributions
As mentioned, our work seeks to understand if 1 is tight, or if it is possible to improve on the existing Weighted MinHash method for communication-free coupling. Towards that end, we study an alternative approach based on the so-called “Gumbel Max-Trick” or “Gumbel sampling”. The method can be described in one sentence: using shared random variables $u_{1},\ldots,u_{n}$ drawn uniformly from $[0,1]$, Alice returns $a=\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{p_{i}}$  [^1] Often $\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{p_{i}}$ is computed as $\operatorname{argmax}_{i\in\{1,\ldots,n\}}\ln p_{i}-\ln(-\ln(u_{i}))$. If $u$ is a uniform random variable in $[0,1]$, then $-\ln(-\ln(u))$ follows a Gumbel distribution (Gumbel, 1935) and hence the name. and Bob returns $b=\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{q_{i}}$, where $p_{i}$ and $q_{i}$ denote the respective probabilities that Alice and Bob’s distributions, $\mathcal{P}$ and $\mathcal{Q}$, assign to item $i$. Intuitively, $a$ and $b$ will be partially coupled because, if some $u_{i}$ happens to be small, we are more likely to set both $a=i$ and $b=i$.
We are interested in Gumbel sampling for two reasons. First, it is already widely used in machine learning for sampling from discrete distributions, including for auto-regressive language generation (Kool et al., 2019). So, using the method for speculative decoding would require essentially no code changes (Maddison et al., 2014; Huijben et al., 2023). Second, the method has been widely studied in theoretical computer science under different names, including PPSWOR and bottom-k sampling (Cohen, 1997; Rosén, 1997; Cohen, 2023). Like Weighted MinHash, Gumbel sampling is a “coordinated random sampling” method, and can be applied to many of the same applications (Bessa et al., 2023). As such it is a natural point of comparison.
We prove the following bound on the performances of Gumbel sampling in Section 3:
Theorem 2.
The communication-free Gumbel coupling method (Protocol 3) samples $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ such that:
$\Pr[a=b]=\sum_{j=1}^{n}\frac{1}{\sum_{i=1}^{n}\max(p_{i}/p_{j},q_{i}/q_{j})}\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}.$
Theorem 2 shows that our protocol based on Gumbel sampling matches the worst-case $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ bound of 1. However, for many distributions, the method often performs much better, suggesting that the inequality in Theorem 2 is can be loose. In particular, in Section 5 we find that the method uniformly outperforms Weighted MinHash for all distributions tested. Proving that Gumbel sampling dominates Weighted MinHash for any pair of distributions is an interesting question for future work.
Nevertheless, despite Gumbel sampling’s good empirical performance, the worst-case bound in Theorem 2 is no better than 1. Our next result shows that this is inherent. In Section 4 we prove:
@@
Theorem 3 establishes that no protocol can improve on Theorem 2 for all values of $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. Perhaps surprisingly, this establishes that extremely simple, time-tested algorithms already yield optimal communication-free coupling (at least when parameterizing the bound via the total variance distance $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$).
Since Theorem 3 essentially rules out major improvements on communication-free coupling, it is natural to ask what is possible if we restrict, but do not entirely eliminate communication. How many bits of communication between Alice and Bob are required to ensure that $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-\epsilon$, i.e., to nearly match what is possible with optimal coupling? A baseline is to simply discretize $\mathcal{P}$ or $\mathcal{Q}$, communicate the entire distribution, and compute a near optimal coupling. This takes $O(n\log(n/\epsilon))$ bits of communication (see Section 6 for details). We show that it is possible to get by with much less. In Section 6 we prove:
@@
There is a protocol that, for any $\epsilon\in(0,1)$, requires $O(\log(n/\epsilon))$ bits of communication in expectation between Alice and Bob to produce samples $a\sim\mathcal{P}$, $b\sim\mathcal{Q}$ such that:
@@
Finally, beyond our theoretical results, we develop our application of communication-free coupling to “drafter-invariant speculative decoding” in Section 5. We provide experimental results which show that the method nearly matches standard speculative decoding in terms of speculation accuracy, but with the added benefit that the neural network output is completely invariant to the drafter network utilized.

### v4

@@ Abstract
Suppose Alice has a distribution $\mathcal{P}$ and Bob has a distribution $\mathcal{Q}$. Alice wants to draw a sample $a\sim\mathcal{P}$ and Bob a sample $b\sim\mathcal{Q}$ such that $a=b$ with as high of probability as possible. It is well-known that, by sampling from an optimal coupling between the distributions, Alice and Bob can achieve $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$, where $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ is the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$.
What if Alice and Bob must solve this same problem without communicating at all? Perhaps surprisingly, with access to public randomness, they can still achieve $\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}\geq 1-2D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ using a simple protocol based on the Weighted MinHash algorithm. This bound was shown to be optimal in the worst-case by Bavarian, Ghazi, Haramaty, Kamath, Rivest, and Sudan [ToC 2020].
In this work, we revisit the “communication-free coupling” problem. We provide a simpler proof of the optimality result from [Bavarian et al., 2020]. Moreover we show that, while the worst-case success probability of Weighted MinHash cannot be improved, an equally simple protocol based on Gumbel sampling offers a Pareto improvement: for every pair of distributions $\mathcal{P}$ and $\mathcal{Q}$, Gumbel sampling achieves an equal or higher value of $\Pr[a=b]$ than Weighted MinHash.
Importantly, this improvement translates to practice. We demonstrate an application of communication-free coupling to speculative decoding, a recent method for accelerating autoregressive large language models [Leviathan, Kalman, Matias, ICML 2023]. We show that communication-free protocols can be used to contruct Drafter-Invariant Speculative Decoding schemes, which have the desirable property that their output is fixed given a fixed random seed, regardless of what drafter is used for speculation. In experiments on a language generation task, Gumbel sampling outperforms Weighted MinHash. Code is available at https://github.com/majid-daliri/DISD.
Finally, we study the coupling problem in the setting where communication is bounded, rather than completely eliminated. We describe a protocol that uses just $O(\log(n/\epsilon))$ bits of communication to achieve $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-\epsilon$, i.e. to essentially match optimal coupling.
We thank Cheuk Ting Li for bringing several related prior works to our attention after we initially published this paper. Angel and Spinka (2021) prove the same result that we do on the Gumbel sampling method, showing that it offers a Pareto improvement over Weighted MinHash. Moreover, in addition to our work and the cited work by Bavarian et al. (2020), Li and Anantharam (2019, 2021) also prove that $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ is optimal in the worst-case for communication-free coupling.
@@ 1 Introduction
The goal of this paper is to revist a “communication-free coupling” problem that arises repeatedly in computer science, and was studied explicitly by Bavarian et al. (2020) under the name correlated sampling. Suppose we have two cooperating parties, Alice and Bob. Alice has a distribution $\mathcal{P}$ and Bob has a distribution $\mathcal{Q}$, which, for simplicity, we assume are discrete distributions over the set $\{1,\ldots,n\}$. The goal is for Alice to sample $a\in\{1,\ldots,n\}$ according to $\mathcal{P}$ and Bob to sample $b\in\{1,\ldots,n\}$ according to $\mathcal{Q}$, in such a way that we maximize the probability that $a=b$. We are interested in how well Alice and Bob can solve this problem without communicating with each other. We do allow them access to a source of shared random numbers.
@@
where $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ is the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$ (Wu, 2020, Section 4.3). To do so, they construct an optimal coupling between the distributions, i.e., a distribution $\mathcal{C}$ over $\{1,\ldots,n\}\times\{1,\ldots,n\}$ with marginals $\mathcal{P}$ and $\mathcal{Q}$ that maximizes $\Pr_{(a,b)\sim\mathcal{C}}[a=b]$. In fact, the total variation distance is defined as $\max_{\text{couplings }\mathcal{C}}\Pr_{(a,b)\sim\mathcal{C}}[a\neq b]$, so (1) is optimal by definition, with or without communication.
While an optimal coupling between $\mathcal{P}$ and $\mathcal{Q}$ has a simple closed form (see Section 2 for details), sampling from the coupling requires knowledge of both $\mathcal{P}$ and $\mathcal{Q}$. I.e., sampling from an optimal coupling requires communication between Alice and Bob. However, it turns out that it is possible to sample from a near-optimal coupling without any communication at all. In particular, the starting point of our work is a straightforward observation from Bavarian et al. (2020), which is implicit in prior work (Manasse et al., 2010):
@@
There is a communication-free protocol (Protocol 2) by which Alice samples $a\sim\mathcal{P}$ and Bob $b\sim\mathcal{Q}$ such that:
@@
Since $0\leq D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})\leq 1$, (2) is always larger than $1-2D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. We conclude that it is possible to almost match the optimum in (1) without communication. 1 follows from an analysis of the popular “Weighted MinHash” method (Kleinberg and Tardos, 2002; Holenstein, 2009; Manasse et al., 2010; Ioffe, 2010; Shrivastava, 2016; Christiani, 2020). We provide a sharp analysis of the method (with an exact expression for $\Pr[a=b]$) in Section 3, 5.
@@ 1.1 Motivation
The main question we address in this work is if it is possible to improve even further on the excellent performance of the existing Weighted MinHash method for communication-free coupling. To cut to the point: it is. We show that a simple method based on the “Gumbel Max-Trick” provides a Pareto improvement over Weighted MinHash. Before discussing this result in detail, however, we provide some motivation for studying communication-free coupling in such depth. It turns out that this problem has a direct application to a technique called speculative decoding that has gained traction for its ability to accelerate inference in autoregressive language models (Leviathan et al., 2023; Chen et al., 2023; Sun et al., 2023), one of the most important algorithmic challenges in modern machine learning (Hoffmann et al., 2022; Zhou et al., 2024).
@@
However, speculative decoding suffers from one major and subtle drawback: if the small “drafter” neural network changes (i.e., because of an update to improve the model), the use of an optimal coupling means that tokens generated by the state-of-the-art neural network will change. This can be potentially problematic in applications of autoregressive language models, where the expected and desired behavior is that the output of a model is fixed given a fixed random seed. This behavior allows for researchers and other users to reliably reproduce results and leads to easier unit testing and debugging.
@@ 1.2 Our Contributions
Motivated by this application, our work seeks to understand if it is possible to improve on the existing Weighted MinHash method for communication-free coupling. Increasing $\Pr[a=b]$ from $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ closer to the optimal $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ could mean up to a 2x reduction in wrong guesses during speculative decoding, which would translate to a potential 2x reduction in computational depth during language generation.
Unfortunately, improving on the bound in 1 is not possible in the worst-case. As shown in an elegant proof by Bavarian et al. (2020), no protocol can do better than $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ in the worst-case. We restate a version [^1] The result in Bavarian et al. (2020) is slightly stronger than our Theorem 2 in that it allows for any total variation distance, not just distances of the form $1/d$ for integer $d$. We assume distance of this form to simplify our argument. of this result below and provide an even more compact proof in Section 3.2.
@@
Importantly, while Theorem 2 rules out an improvement on Weighted MinHash for all distribution pairs, it does not mean that the method is optimal. Our main contribution is to demonstrate that another extremely simple protocol offers a Pareto improvement over Weighted MinHash: for any $\mathcal{P},\mathcal{Q}$ it performs at least as well, but sometimes performs better (often significantly so in practice). Our alternative approach is based on the so-called “Gumbel Max-Trick” or “Gumbel sampling”. The method can be described in one sentence: using shared random variables $u_{1},\ldots,u_{n}$ drawn uniformly from $[0,1]$, Alice returns $a=\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{p_{i}}$  [^2] Often $\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{p_{i}}$ is computed as $\operatorname{argmax}_{i\in\{1,\ldots,n\}}\ln p_{i}-\ln(-\ln(u_{i}))$. If $u$ is a uniform random variable in $[0,1]$, then $\ln(-\ln(u))$ follows a Gumbel distribution (Gumbel, 1935) and hence the name. and Bob returns $b=\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{q_{i}}$, where $p_{i}$ and $q_{i}$ denote the respective probabilities that Alice and Bob’s distributions, $\mathcal{P}$ and $\mathcal{Q}$, assign to item $i$. Intuitively, $a$ and $b$ will be partially coupled because, if some $u_{i}$ happens to be small, we are more likely to set both $a=i$ and $b=i$.
Gumbel sampling has been widely studied in theoretical computer science under different names, including PPSWOR and bottom-k sampling (Cohen, 1997; Rosén, 1997; Cohen, 2023). Moreover, it is already widely used in machine learning for uncoupled sampling from discrete distributions, including for auto-regressive language generation (Kool et al., 2019). So, using the method for speculative decoding would require essentially no code changes (Maddison et al., 2014; Huijben et al., 2023). We prove the following bound on the performances of Gumbel sampling in Section 3:
@@
For distributions $\mathcal{P}$ and $\mathcal{Q}$, let $C_{\text{WMH}}(\mathcal{P},\mathcal{Q})$ be the probability that samples $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ drawn using the Weighted MinHash method (Protocol 2) satisfy $a=b$. The communication-free Gumbel coupling method (Protocol 3) samples $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ such that:
$\Pr[a=b]=\sum_{\begin{subarray}{c}j\in\{1,\ldots,n\}\\ \min(p_{j},q_{j})>0\end{subarray}}\frac{1}{\sum_{i=1}^{n}\max(p_{i}/p_{j},q_{i}/q_{j})}\geq C_{\text{WMH}}(\mathcal{P},\mathcal{Q})\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}.$
It is not hard to find distributions for which the first inequality above is strict, i.e., for which Gumbel sampling strictly improves on Weighted MinHash. In fact, as discussed in Section 3, this will always be the case under the mild condition that $p_{i}\neq q_{i}$ for at least three values of $i$. As an example, consider distributions $\mathcal{P}=\{{1}/{2},{1}/{2},0\}$ and $\mathcal{Q}=\{{1}/{3},{1}/{3},{1}/{3}\}$ over $\{1,2,3\}$. The best possible collision probability for these distributions is $1-\mathcal{D}_{TV}(\mathcal{P},\mathcal{Q})={2}/{3}$ and $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}=1/2$. In this case, Gumbel sampling actually obtains the optimal collision probability of 2/3{2}/{3}, while Weighted MinHash obtains probability $7/12$, closer to the worst-case bound for communication-free methods. A similar pattern emerges in our experiments with real-word next-token distributions that arise in language generation Section 5. Gumbel sampling often nearly matches the optimal coupling, obtaining collision probability close to $1-\mathcal{D}_{TV}(\mathcal{P},\mathcal{Q})$, while Weighted MinHash performs somewhat worse.
It is interesting to ask if additional improvement is possible. Is there another protocol that offers a Pareto improvement over Gumbel sampling, or is the approach Pareto optimal? I.e., does improvement for a pair of distributions $(\mathcal{P},\mathcal{Q})$ necessarily require worse performance for another pair?
Considering the worst-case bound of Theorem 2, it is also natural to ask what is possible if we restrict, but do not entirely eliminate communication. How many bits of communication between Alice and Bob are required to ensure that $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-\epsilon$, i.e., to nearly match what is possible with optimal coupling? A baseline is to simply discretize $\mathcal{P}$ or $\mathcal{Q}$, communicate the entire distribution, and compute a near optimal coupling. This takes $O(n\log(n/\epsilon))$ bits of communication (see Section 6 for details). We show that it is possible to get by with much less. In Section 6 we prove:
Theorem 4.
There is a protocol that, for any $\epsilon\in(0,1)$, requires a constant number of rounds and $O(\log(n/\epsilon))$ bits of communication, in expectation, between Alice and Bob to produce samples $a\sim\mathcal{P}$, $b\sim\mathcal{Q}$ such that:
@@
Finally, beyond our theoretical results, we implement our Gumbel sampling method and evaluate its performance in the “drafter-invariant speculative decoding” task in Section 5. We show that the method nearly matches standard speculative decoding in terms of speculation accuracy, but with the added benefit that the neural network output is completely invariant to the drafter network utilized.